A rocket is launched from ground level with an initial velocity of 224 ft/s. The height (h) in feet of the rocket at any given time h(t) =224t-16t^2
a: what does (8768) mean in the context of the problem?
b: when will the rocket reach height of 528 ft?
c:when will the rocket reach ground?
D: when will the rocket reach its maximum height?
e: what is the maximum height of the rocket?
f: what is the domain and range for the SITUATION?

Answer by richwmiller(17219)

h (t) =224t-16t^2
a) no clue what 8768 means
16*548=8768
2^6×137=8768
b)
528 =224t-16t^2
at t=3 and t=11
c)
0 =224t-16t^2
at t=0 and at t=14
d)
h(t) =-16t^2+224t
-b/2a
224/32=7 at t =7 it reaches its maximum height
e) max height is 784 at t=7
f) domain : all real numbers
range : real numbers where y<=784}